Reese, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late December, she received a $20,000 bill from her accountant for consulting services related to her small business. Reese can pay the $20,000 bill anytime before January 30 of next year without penalty. Assume Reese’s marginal tax rate is 32 percent this year and 35 percent next year, and that she can earn an after-tax rate of return of 12 percent on her investments.

a. What is the after-tax cost if she pays the $20,000 bill in December?

b. What is the after-tax cost if she pays the $20,000 bill in January?

c. Should Reese pay the $20,000 bill in December or January?

d. What is the after-tax cost of she expect her marginal tax rate to be 24 percent next year and pays the $20,000 bill in January?

e. Should Reese pay the $20,000 bill in December or January if she expects her marginal tax rate to be 32 percent this year and 24 percent next year?
Answer #1

a) 20000*32% = 6400

20000 - 6400 = $13,600

If she pays in December then the after tax cost = $ 13600

b) 20000*35% = 7000

Present Value of tax = $7000* 0.893( discount factor, 1 year, 12%)

Present value of tax = $6251

After tax cost = Pre-tax cost - present value tax

After tax cost = 20000 - 6251

After tax cost of January = $13,749

c) Since the after tax cost of $13600 is lower then $13749 then, Reese should pay the bill in December of $13600.

d) If she payes her bill in January with 24% tax rate then

20000*24%= 4800

Present value of tax = $4800* 0.893 (discount factor, 1 year, 12%)

Present value of tax = $4286.4

After tax cost = Pre-tax cost - present value tax

After tax cost= 20000 - 4286.4

After tax cost of January = $ 15713.6

e) In this case also Reese cost is lower in December month.

So she should pay $13600 in December.